### v1

@@ 1 INTRODUCTION
Few-shot learning [8, 9, 10] can enhance music understanding by addressing two key challenges. First, unlike certain transfer learning methods that are limited to frequent classes [11, 12, 13], it enables the inclusion of under-represented tags within well-established music domains, where annotated data is often sparse. Second, it supports the exploration of under-represented domains, such as new or emerging music cultures, where annotated datasets are even more limited.
In order to effectively address the automatic audio tagging problem, the adoption of multi-label few-shot learning (ML-FSL) methods is essential. While these methods have been employed in recent audio classification research [14], to the best of our knowledge, they have not yet been applied to MIR datasets. Such a technique can, in turn, be combined with either foundation models [15] or contrastive learning models [16, 17, 18], mapping their general-purpose representations to semantic spaces for more comprehensive music analysis.
@@ 2.1 AUTOMATIC WORLD MUSIC TAGGING
In order to describe a piece of music with metadata, various aspects can be considered, including genre, mood, and instrumentation. It is a common approach in the MIR field to regard this non-exhaustive list of metadata as tags that characterize a music work [19]. The task of predicting the tags of a music piece given its audio signal is referred to as auto-tagging [4, 5, 6, 7], constituting a multi-label classification problem, where an instance can be annotated with a set of multiple labels.
@@ 3.2 LC-PROTONETS
Figure 1: The labels of the support set consisting of four items (bottom) and the LC-classes created from them (top). The concentric circles with the dashed lines pass through LC-Protonets with same representations in the embedding space that are graphed equidistant from the query item $q$.
@@
Figure 1 illustrates an indicative example. The support set $S$ consists of four items, each associated with a distinct set of labels. The set of label combinations, henceforth referred to as LC-classes, is defined as the union of the power sets of each $\mathbf{y}_{i}$ in the support set. Each LC-class is represented by an LC-Protonet (LCP), whose representation is computed by averaging the embeddings of the support items that include the corresponding LC-class in their power set, as shown in the color-coded example in the figure.
@@
For a query item $q\in Q$, the distances to all LC-Protonets are computed. Figure 1 provides a conceptual 2D representation of the embedding space, where concentric circles indicate equal distances from $q$ to different LCPs. In the actual space, the LC-Protonets with identical representations will occupy the same point.
In cases where a query item has equal distances to multiple LC-Protonets, the one representing the largest number of labels is selected. That way, hierarchical relationships and strong correlation between the labels are supported. In the depicted example, both the $\{E\}$ and $\{C,E\}$ LCPs have the minimum distance to $q$, and the query is assigned to the $\{C,E\}$ LC-class. If $E$ is hierarchically subordinate to $C$, the LCPs for $\{E\}$ and $\{C,E\}$ will share the same representation, but the model will consistently select the $\{C,E\}$ class. Even if $C$ and $E$ are not part of a formal hierarchy but they co-occur in the same support items, it is rational to assume strong correlation between them and, again, assign the query to the $\{C,E\}$ class.
@@
The LC-Protonet representation $\mathbf{p}_{j}$ for the corresponding class is computed by averaging the embeddings of the items in $S_{j}$:
@@
In each episode, a specified number of query items for each active label is sampled to form the query set $Q$. Given a query item $\mathbf{x}_{i}$ with an initial label vector $\mathbf{y}_{i}\in\{0,1\}^{N}$ (a multi-hot vector of dimension $N$), the expanded label vector $\mathbf{y}_{mH_{i}}\in\{0,1\}^{|L|}$ is constructed by assigning a value of 1 to each of the item’s LC-classes: $\mathbf{y}_{mH_{i}}(j)=1\quad\text{if}\quad L_{j}\in\mathcal{P}(\mathbf{y}_{i}),\quad\forall j=1,2,...,|L|.$ The loss function is based on the binary cross-entropy:
$\begin{split}Loss(\mathbf{x}_{i},&\mathbf{y}_{mH_{i}},\mathbf{p})=-\mathbf{y}_{mH_{i}}\log(\sigma(-d(f_{\theta}(\mathbf{x}_{i}),\mathbf{p})))+\\ &(1-\mathbf{y}_{mH_{i}})\log(1-\sigma(-d(f_{\theta}(\mathbf{x}_{i}),\mathbf{p}))),\end{split}$ (7)
where $d$ is the distance function, $\sigma$ the sigmoid function and $\mathbf{p}$ the LC-Protonets representations. We minimize the loss for all items in the query set $Q$.
Inference phase: At the inference phase, the support set $S$ is created following a similar “N-way K-shot” setup used during training. The LC-classes $L$ and their corresponding LC-Protonet representations $\mathbf{p}_{j}$ (for $j=1,2,...,|L|$) are computed. For a query instance, the distances to all LC-Protonets are calculated, and the instance is assigned to the LC-class represented by the nearest LC-Protonet:
@@ 4 EXPERIMENTS
Table 2: Macro-F1 (M-F1) and micro-F1 (m-F1) scores (%) with standard deviation, when aggregating the performance of the three methods over all datasets and training setups, for various “N-way” ML-FSL tasks, with a “3-shot” approach applied to all of them.
@@
ML-PNs 65.21±3.48 66.12±3.55 46.32±3.12 45.92±2.85 39.31±1.34 44.23±1.7 21.45±1.05 21.02±0.99
One-vs.-Rest 64.69±3.37 65.84±3.35 42.69±2.85 42.5±2.74 35.44±1.32 39.4±1.44 18.8±1.17 18.56±1.13
LC-Protonets (ours) 62.6±3.97 66.26±3.92 47.89±5.39 49.34±4.98 42.84±2.19 56.28±2.3 28.5±2.91 31.37±3.02
@@
ML-PNs 29.84±0.92 32.57±0.97 24.61±0.66 29.26±1.02 19.74±0.53 23.05±0.82 17.49±0.49 19.45±0.54
One-vs.-Rest 25.64±0.97 27.81±1.09 21.74±0.79 25.36±1.04 17.06±0.61 19.58±0.8 14.79±0.55 16.17±0.69
LC-Protonets 36.77±1.97 50.77±1.44 31.31±1.73 52.65±1.57 28.09±1.42 50.28±1.65 28.45±1.51 46.48±1.47
4.1 DATASETS AND METRICS
@@
For evaluation, we selected Macro-F1 and Micro-F1 scores. Since our method directly predicts a set of labels without assigning probabilities to individual labels, calculating metrics like Area Under the Curve (AUC) is not straightforward. F1 score is the harmonic mean of the precision and recall scores and in its Macro- setting, the mean of the individual label scores is calculated. Micro-F1 computes metrics globally by aggregating true positives, false negatives, and false positives across all samples.
@@ 4.4 EXPERIMENTAL SETUP
We split the labels of each dataset into training and testing sets. The training set is used both during the training phase of ML-FSL models and for pre-training the backbone VGG-ish model via supervised learning, while the testing set is used to form the novel classes for evaluation. For the audio recordings, we use the same splits for training, validation, and testing as in previous studies [11, 6].
To train the ML-FSL models, we employ three setups: (i) training from scratch with random weight initialization, (ii) full fine-tuning of a pre-trained backbone model, and (iii) fine-tuning only the last layer of the pre-trained backbone model. The model architecture remains the same across all setups, except that in the two fine-tuning cases, a VGG-ish model pre-trained on the well-represented tags is transferred as the backbone, with only the final classification layer excluded.
@@
Cosine distance is used as the distance metric, and all methods are trained using episodic learning with a $12$-way $3$-shot setup. This setup is chosen to accommodate the low-resource nature of music data, as the absolute number of labels differs across domains. For instance, Hindustani, Carnatic, and FMA-medium datasets have only 20 labels in the training split. Additionally, selecting 3 examples per label allows under-represented labels to be included. $50$ episodes are sampled for each epoch and $3$ query items per label are utilized to compute the loss in each one of them.
The validation set is formed by holding out 5 classes from the training set during learning. The Adam optimizer [54] is used, and early stopping is applied based on the Macro-F1 score on the validation set at the end of each epoch. Regarding the input, a 512-point FFT with a 50% overlap is applied to the audio signal, and the Mel bands are set to 128. During training, a random chunk of each audio recording is selected, while during testing, the average embedding of all chunks forms the representation of an instance.
@@ 4.5 TWO-STEP LEARNING METHOD
In imbalanced datasets, it is common to encounter a large number of labels that occur infrequently, leading to a long-tailed label distribution. When training models using supervised learning, a threshold is often set to include only the most frequent categories. To address this limitation, we propose a two-step method that combines supervised and few-shot learning. Unlike the ML-FSL training setups that use a pre-trained backbone model, no fine-tuning is applied in this approach.
@@
Table 3: Macro-F1 (M-F1) and micro-F1 (m-F1) scores (%) for a “30-way 3-shot” task with “Base & Novel” classes, under four conditions: (i) training from scratch, (ii) full fine-tuning of a pre-trained backbone model, (iii) fine-tuning of the final layer of a pre-trained backbone, and (iv) using a pre-trained backbone without any fine-tuning. Rows represent the multi-label few-shot learning methods, while columns correspond to the different datasets.
@@ 5.1 ML-FSL TASKS
In Table 2, the aggregated results of the LC-Protonets method and the two comparative approaches are presented. These results were calculated by averaging performance across all datasets and training setups: from scratch, full fine-tuning, and fine-tuning of the last layer. Each experiment was run five times, and the standard deviation is reported. We present both macro-F1 and micro-F1 scores for different numbers of labels, ranging from $5$ to $60$. The evaluations were performed on “Base” and “Novel” classes for smaller numbers of classes, and on “Base & Novel” classes for larger numbers.
@@
In terms of standard deviation, we observed larger values for fewer-shot conditions, such as the $5$-way task, due to the random sampling of a small number of active classes in imbalanced datasets. Additionally, LC-Protonets’ reliance on support set sampling for deriving LC-classes leads to higher standard deviation values compared to the other methods.
Figure 2 highlights the differences between the prototype representations generated by ML-PNs and LC-Protonets. While ML-PNs create one prototype per class, LC-Protonets populate the embedding space with representations derived from the power sets of the support item labels. We believe this enhanced positive sampling of the feature space contributes to the significant performance improvement seen in the results.
@@
Finally, the last three rows of Table 3 report the performance when a pre-trained backbone model was used without any fine-tuning. Interestingly, LC-Protonets maintained performance levels similar to those seen in the fine-tuning setups. This suggests that the method relies more on the quality of the representations provided by the backbone model than on episodic learning. By contrast, the performance of the comparative approaches dropped significantly, as they encountered challenges with multi-label classification without training, often assigning all labels to all samples. This can be seen in the zero standard deviation of ML-PNs on the FMA-medium dataset, for example.
Figure 4: Scalability metrics of the proposed method, averaged across all datasets. The $x$-axis represents the number of labels, the left $y$-axis shows the number of LC-Protonets, and the right $y$-axis indicates the inference time per item. Both $y$-axes use the same logarithmic scale.
Table 4: Macro-F1 (M-F1) and micro-F1 (m-F1) scores (%) for the “VGG-ish” model on the set of well-represented tags, and for the two-step method “VGG-ish & LCPs” (VGG-ish & LC-Protonets) on both the well-represented and extended tag sets for each dataset.
@@ 5.2 TWO-STEP LEARNING METHOD
The results of the proposed two-step learning method are shown in Table 4. For each dataset, two different numbers of tags are used. The smaller number, such as $20$ for FMA-medium, corresponds to the well-represented tags on which the VGG-ish model was trained, while the larger number, $40$, represents the extended set of tags. The performance of both the “VGG-ish” model and the “VGG-ish & LCPs” (VGG-ish & LC-Protonets) method on the well-represented tags is reported, and for the latter, its performance on the extended tag set is also included.
@@
The percentage drop in macro-F1 performance is significantly smaller than the percentage increase in the number of tags. For example, in the Turkish-makam dataset, the macro-F1 score drops from $37.23\%$ to $30.07\%$ as the number of tags increases from $30$ to $60$. An exception is the Lyra dataset, where performance on the extended tag set remains nearly identical to the well-represented tags. This could be due to stronger correlations between tags in Lyra compared to the other datasets.

### v2

@@ 1 INTRODUCTION
In our previous work [8], we utilized transfer learning methods to enhance music understanding with a cross-cultural perspective. However, like similar approaches [9, 10], this method is limited to frequent classes. Traditional multi-label classification methods, such as [11], which model label correlations by constructing random subsets of labels, are effective but not designed for scenarios where data is minimal and labels are often imbalanced. Few-shot learning [12, 13, 14] addresses these limitations by enabling the inclusion of under-represented tags within well-established music domains, where annotated data is often sparse. Furthermore, it supports the exploration of under-represented domains, such as new or emerging music cultures, where annotated datasets are even more limited.
In order to effectively address the automatic audio tagging problem, the adoption of multi-label few-shot learning (ML-FSL) methods is essential. While ML-FSL methods have been employed in recent audio classification research [15], to the best of our knowledge, they have not yet been applied to MIR datasets. Such a technique can, in turn, be combined with either foundation models [16] or contrastive learning models [17], mapping their general-purpose representations to semantic spaces for more comprehensive music analysis.
@@ 2.1 AUTOMATIC WORLD MUSIC TAGGING
In order to describe a piece of music with metadata, various aspects can be considered, including genre, mood, and instrumentation. It is a common approach in the MIR field to regard this non-exhaustive list of metadata as tags that characterize a music work [18]. The task of predicting the tags of a music piece given its audio signal is referred to as music auto-tagging [4, 5, 6, 7], constituting a multi-label classification problem, where an instance can be annotated with a set of multiple labels.
@@ 3.2 LC-PROTONETS
Figure 1: The labels of the support set consisting of four items (bottom) and the LC-classes created from them (top). The concentric circles with the dashed lines pass through LCPs with same representations in the embedding space that are graphed equidistant from the query item $q$.
@@
Figure 1 illustrates an indicative example. The support set $S$ consists of four items, each associated with a distinct set of labels. The set of label combinations, henceforth referred to as LC-classes, is defined as the union of the power sets of each $\mathbf{y}_{i}$ in the support set. Each LC-class is represented by an LC-Prototype (LCP), whose representation is computed by averaging the embeddings of the support items that include the corresponding LC-class in their power set, as shown in the color-coded example in the figure.
@@
For a query item $q\in Q$, the distances to all LCPs are computed. Figure 1 provides a conceptual 2D representation of the embedding space, where concentric circles indicate equal distances from $q$ to different LCPs. In the actual space, the LCPs with identical representations will occupy the same point.
In cases where a query item has equal distances to multiple LCPs, the one representing the largest number of labels is selected. That way, hierarchical relationships and strong correlation between the labels are supported. In the depicted example, both the $\{E\}$ and $\{C,E\}$ LCPs have the minimum distance to $q$, and the query is assigned to the $\{C,E\}$ LC-class. If $E$ is hierarchically subordinate to $C$, the LCPs for $\{E\}$ and $\{C,E\}$ will share the same representation, but the model will consistently select the $\{C,E\}$ class. Even if $C$ and $E$ are not part of a formal hierarchy but they co-occur in the same support items, it is rational to assume strong correlation between them and, again, assign the query to the $\{C,E\}$ class.
@@
The LCP representation $\mathbf{p}_{j}$ for the corresponding class is computed by averaging the embeddings of the items in $S_{j}$:
@@
In each episode, a specified number of query items for each active label is sampled to form the query set $Q$.
Given a query item $\mathbf{x}_{i}$ with a label set $\mathbf{y}_{i}$, its initial multi-hot label vector $\mathbf{z}_{i}\in\{0,1\}^{N}$ is constructed such that $\mathbf{z}_{i}(k)=1\quad\text{if}\quad k\in\mathbf{y}_{i},\forall k=1,2,\ldots,N$. The expanded multi-hot vector $\mathbf{z}_{mH_{i}}\in\{0,1\}^{|L|}$ is then constructed by assigning a value of 1 to each of the item’s LC-classes: $\mathbf{z}_{mH_{i}}(j)=1\quad\text{if}\quad L_{j}\in\mathcal{P}(\mathbf{y}_{i}),\quad\forall j=1,2,...,|L|.$ The loss function is based on the binary cross-entropy:
$\begin{split}Loss(\mathbf{x}_{i},&\mathbf{z}_{mH_{i}},\mathbf{p})=-\mathbf{z}_{mH_{i}}\log(\sigma(-d(f_{\theta}(\mathbf{x}_{i}),\mathbf{p})))+\\ &(1-\mathbf{z}_{mH_{i}})\log(1-\sigma(-d(f_{\theta}(\mathbf{x}_{i}),\mathbf{p}))),\end{split}$ (7)
where $d$ is the distance function, $\sigma$ the sigmoid function and $\mathbf{p}$ the LCPs representations. We minimize the loss for all items in the query set $Q$.
Inference phase: At the inference phase, the support set $S$ is created following a similar “N-way K-shot” setup used during training. The LC-classes $L$ and their corresponding LCPs representations $\mathbf{p}_{j}$ (for $j=1,2,...,|L|$) are computed. For a query instance, the distances to all LCPs are calculated, and the instance is assigned to the LC-class represented by the nearest LCP:
@@
4 EXPERIMENTS
4.1 DATASETS AND METRICS
Table 2: Macro-F1 (M-F1) and micro-F1 (m-F1) scores (%) with 95% confidence intervals, when aggregating the performance of the three methods over all datasets and training setups, for various “N-way” ML-FSL tasks, with a “3-shot” approach applied to all of them.
@@ 4.1 DATASETS AND METRICS
ML-PNs 65.21±4.33 66.12±4.41 46.32±3.88 45.92±3.53 39.31±1.66 44.23±2.11 21.45±1.3 21.02±1.22
One-vs.-Rest 64.69±4.19 65.84±4.16 42.69±3.53 42.5±3.4 35.44±1.64 39.4±1.79 18.8±1.45 18.56±1.41
LC-Protonets (ours) 62.6±4.93 66.26±4.87 47.89±6.69 49.34±6.19 42.84±2.71 56.28±2.86 28.5±3.61 31.37±3.74
@@
ML-PNs 29.84±1.14 32.57±1.2 24.61±0.83 29.26±1.27 19.74±0.66 23.05±1.02 17.49±0.62 19.45±0.67
One-vs.-Rest 25.64±1.21 27.81±1.35 21.74±0.98 25.36±1.29 17.06±0.76 19.58±1.0 14.79±0.68 16.17±0.86
LC-Protonets 36.77±2.44 50.77±1.79 31.31±2.14 52.65±1.95 28.09±1.76 50.28±2.06 28.45±1.87 46.48±1.82
@@
Since our method directly predicts a set of labels without assigning probabilities to individual labels, calculating metrics like Area Under the Curve (AUC) is not straightforward. Therefore, we selected Macro-F1 and Micro-F1 scores for evaluation. F1 score is the harmonic mean of the precision and recall scores and in its Macro- setting, the mean of the individual label scores is calculated. Micro-F1 computes metrics globally by aggregating true positives, false negatives, and false positives across all samples.
@@ 4.4 EXPERIMENTAL SETUP
We split the labels of each dataset into training and testing sets. The training set is used both during the training phase of ML-FSL models and for pre-training the backbone VGG-ish model via supervised learning, while the testing set is used to form the novel classes for evaluation. For the audio recordings, we use the same splits as in previous studies [8, 6], with a split ratio of $0.7$, $0.1$, and $0.2$ for the training, validation, and test sets, respectively.
To train the ML-FSL models, we employ three setups: (i) training from scratch with random weight initialization, (ii) full fine-tuning of a pre-trained backbone model, and (iii) fine-tuning only the last layer of the pre-trained backbone model. The model architecture remains the same across all setups, except in the fine-tuning cases where a VGG-ish model pre-trained on the well-represented tags is transferred as the backbone, excluding only the final classification layer.
@@
Cosine distance is used as the distance metric, and all methods are trained using episodic learning with a $10$-way $3$-shot setup. This setup is chosen to accommodate the low-resource nature of music data, as the absolute number of labels differs across domains. For instance, Hindustani, Carnatic, and FMA-medium datasets have only 20 labels in the training split. Additionally, selecting 3 examples per label allows under-represented labels to be included. $50$ episodes are sampled for each epoch and $3$ query items per label are utilized to compute the loss in each one of them.
The validation set is formed by holding out 5 classes from the training set during learning. The Adam optimizer [53] is used, and early stopping is applied based on the Macro-F1 score on the validation set. Regarding the input, the audio signal is sampled at 16 kHz, and a 512-point FFT with a 50% overlap is applied while the Mel bands are set to 128. During training, a random chunk of each audio recording is selected, while during testing, the average embedding of all chunks forms the representation of an instance.
@@ 4.5 TWO-STEP LEARNING METHOD
In imbalanced datasets, it is common to encounter a large number of labels that occur infrequently, leading to a long-tailed label distribution. When training models using supervised learning, a threshold is often set to include only the most frequent categories. To address this limitation, we propose a two-step method that combines supervised and few-shot learning. Unlike ML-FSL setups that fine-tune a pre-trained backbone, this approach requires no fine-tuning.
@@
Table 3: Macro-F1 (M-F1) and micro-F1 (m-F1) scores (%) with 95% confidence intervals for a “30-way 3-shot” task with “Base & Novel” classes under four conditions: (i) training from scratch, (ii) full fine-tuning of a pre-trained backbone, (iii) fine-tuning only the final layer, and (iv) using a pre-trained backbone without fine-tuning. Rows represent the multi-label few-shot learning methods, and columns correspond to the datasets.
@@ 5.1 ML-FSL TASKS
In Table 2, the aggregated results of the LC-Protonets method and the two comparative approaches are presented. These results were calculated by averaging performance across all datasets and training setups: from scratch, full fine-tuning, and fine-tuning of the last layer. Each experiment was run five times with different random seeds, and the $95\%$ confidence intervals are reported. While label splits remained consistent across runs, different active classes were sampled at each epoch during training, and different support items were selected in each run. We present both macro-F1 and micro-F1 scores for different numbers of labels, ranging from $5$ to $60$. The evaluations were performed on “Base” and “Novel” classes for smaller numbers of classes, and on “Base & Novel” classes for larger numbers.
@@
In terms of confidence intervals, we observed wider ranges for few-shot conditions, such as the $5$-way task, due to the random sampling of a small number of active classes in imbalanced datasets. Additionally, LC-Protonets’ reliance on support set sampling for deriving LC-classes leads to wider confidence intervals compared to the other methods.
Figure 3 highlights the differences between the prototypes formed by ML-PNs and LC-Protonets. While ML-PNs create one prototype per class, LC-Protonets populate the embedding space with representations derived from the power sets of the support item labels. We believe this enhanced positive sampling of the feature space contributes to the significant performance improvement seen in the results.
@@
Figure 4: Scalability metrics of the proposed method, averaged across all datasets. The $x$-axis represents the number of labels, the left $y$-axis shows the number of LC-Prototypes, and the right $y$-axis indicates the inference time per item. Both $y$-axes use the same logarithmic scale.
@@
Finally, the last three rows of Table 3 report the performance when a pre-trained backbone model was used without any fine-tuning. Interestingly, LC-Protonets maintained performance levels similar to those seen in the fine-tuning setups. This suggests that the method relies more on the quality of the representations provided by the backbone model than on episodic learning. By contrast, the performance of the comparative approaches dropped significantly, as they encountered challenges with multi-label classification without training, often assigning all labels to all samples. This can be seen in the very narrow confidence intervals of ML-PNs on the FMA-medium dataset, for example.
Comparisons with state-of-the-art methods are not possible due to the lack of ML-FSL results for MIR and these datasets. Moreover, the literature commonly reports ROC-AUC for multi-label tasks instead of Macro-F1. State-of-the-art models also focus on top-N labels, unlike our method which targets under-represented classes.
Table 4: Macro-F1 (M-F1) and micro-F1 (m-F1) scores (%) with 95% confidence intervals, for the “VGG-ish” model on the set of well-represented tags, and for the two-step method “VGG-ish & LC-Protonets” on both the well-represented and extended tag sets for each dataset.
@@ 5.2 TWO-STEP LEARNING METHOD
The results of the proposed two-step learning method are shown in Table 4. For each dataset, two tag counts are used. The smaller number, such as $20$ for FMA-medium, corresponds to the well-represented tags on which the VGG-ish model was trained, while the larger number, $40$, represents the extended tag set. The performance of both the “VGG-ish” model and the “VGG-ish & LC-Protonets” method on the well-represented tags is reported, and for the latter, its performance on the extended tag set is also included.
@@
There is a relatively small drop in macro-F1 performance as the number of tags increases significantly. For example, in the Turkish-makam dataset, the macro-F1 score drops from $37.23\%$ to $30.07\%$ as the number of tags rises from $30$ to $60$. An exception is the Lyra dataset, where performance on the extended tag set remains nearly identical to the well-represented tags, likely due to stronger correlations between tags in Lyra compared to the other datasets.